# Global Streams, Local Currents: A Data Analysis on Global VOD Content Consumption

### v1

@@ 3.2 The role of key content-producing countries in similar consumption grouping
Figure 3: Similar consumption groups by the leading content-producing countries. (a)Similar consumption groups by US-produced content, (b)Similar consumption groups by South Korea-produced content, (c)Similar consumption groups by Columbia-produced content
Then, to what extent do the key content-producing countries influence the grouping into similar consumption clusters? To understand how much influence the significant content-producing countries have on the formation of consumption clusters, we conducted the Louvain algorithm process again, but this time we only included the consumption data from the leading content-producing country within each of the three identified consumption groups. Networks were constructed from 528 pieces of United States content, 203 pieces of South Korean content, and 27 pieces of Colombian content.
When forming similar consumption clusters by the leading contries’ content, all countries, except Turkiye, always belonged to the same group in 1,000 runs of the Louvain algorithm. Similar to the previous results, most countries consistently belonged to the same group in 1,000 runs of the Louvain algorithm, but some countries were classified into several groups. Nevertheless, all countries belonged to a certain group at least 700 times out of 1,000, and the groups that countries mainly belonged to were represented in Fig. 3.
@@
Finally, when the content consumption patterns are analyzed with Colombian content as the focus, we observed two main cluster groups: a Hispanic Latin America and Spain group and other remaining groups. That is, when looking from the perspective of Colombia, the division of groups appears to be heavily influenced by language, specifically whether the countries are Spanish-speaking or not. This reflects the cultural and linguistic influence on media consumption within the VOD market.
@@
In sum, US content is similarly consumed in Asia, Central and South America, as well as the Middle East and Africa. Specifically, Spain is classified within the North America and Pan-Europe consumption group when it comes to US content, while Turkiye is classified within the Middle East and Africa. This implies that audiences in Spain and Turkiye consume US content in ways that align with the consumption pattenrs of North America and Pan-Europe, and the Middle East and Africa, respectively. South Korean content consumption closely resembles the three similar consumption clusters previously identified. However, countries like Argentina, Spain, and Turkiye appear to consume South Korean content from a North American and Pan-European perspective. Finally, Colombian content is predominantly consumed within the Spanish-speaking cultural sphere, including Latin America’s central western regions and Spain in Europe. All other countries are seen as part of a different cultural sphere. Overall, this analysis suggests that major content-producing countries influence global content consumption patterns, which can align or differ based on cultural and linguistic ties. The way content from these key countries is consumed can help to define or redefine cultural spheres and clusters within the VOD space.
3.3 Influencing trend of content-producing countries
@@
Figure 4: Trends in the rankings of US, South Korean, and Colombian content by similar consumption groups. The red triangle represents the scores for American content, the green circle represents South Korean content, and the blue star represents Colombian content. (a)The Asia and Middle East Group, (b)The North America and Pan-Europe Group, (c)The Central and South America Group
@@
Based on the original analysis, which defined three similar consumption groups by considering all content, we further examined how the content rankings of leading countries within the VOD similar consumption groups, such as the US, South Korea, and Coloumbia, have changed over the data period. This is to determine whether the previously observed content performance of these countries is due to random variation or if it represents a significant and sustained influence. Figure 4 represents graphs that illustrate the quarterly ranking scores of content from the United States, South Korea, and Colombia within the three established similar consumption groups.
In the Asia and Middle East group, there is a close competition in the preference for content from the US and South Korea, but after the third quarter of 2021, when ‘Squid Game’ was released, the preference for South Korean content has consistently increased. In the North America and Pan-Europe group, while American content is significantly preferred, there is a spike in interest for South Korean content around the time ‘Squid Game’ was released, but it soon returns to its original level. In the Central and South America group, South Korean content consumption, which was previously minimal, increases following the appearance of ‘Squid Game’ and maintains a certain level thereafter.
As a result, it can be seen that the global response to South Korean content has changed significantly since the global creative effort of VOD intensified, particularly after 2021 with ‘Squid Game’ as a turning point. In Asia, South Korean content surpasses the consumption of American content, and it shows meaningful growth in the Central and South American region. The temporary reaction in the North America and Pan-Europe group suggests that the cultural force of these groups is strong.
4 Summary and Discussion
4.1 Summary
@@
Our findings show that global countries are categorized into three groups with geographic and linguistic connections based on global VOD consumption distance: the ‘North America and Pan-Europe group,’ the ‘Asia and Middle East group,’ and the ‘Central and South America group.’ In these three groups, the ‘North America and Pan-Europe group’ is most influenced by the United States, the ‘Asia and Middle East group’ by South Korea, and the ‘Central and South America group’ by the United States and Colombia.
Looking specifically at the top 5 countries, the United States is the dominant primary driving nation in all three groups, followed by Spain and South Korea exerting a mid-range level of influence, and other countries standing out based on their economic and cultural power within the respective groups. This partially aligns with the fact that Netflix’s original investment in Europe is made discriminately according to the existing cultural power[2]. It allows us to consider the cultural power of the platform-dominant United States, the distinctive geographical and linguistic expansion of Spain, and the status of so-called culturally advanced countries within each similar group. Exceptionally, South Korea is a special case that cannot be explained geographically or linguistically within this explanatory framework, warranting further investigation.
From the results of content consumption patterns over time, we observed the impact of a highly popular South Korean series, ‘Squid Game’. After ‘Squid Game’, the preference for South Korean content consistently increased in the ‘Asia and Middle East group’ and the ‘Central and South America group’, while in the ‘North America and Pan-Europe group’, the preference for South Korean content quickly subsided due to the overwhelming influence of the United States. These findings indicate that the release of culturally significant content such as ‘Squid Game’ has the potential to reshape global content consumption patterns, even in regions with traditionally dominant preferences. South Korean content, riding on the wave of ‘Squid Game’s’ unprecedented international success, has notably shifted its global standing and consumer appeal, surpassing American content in Asia and marking significant inroads in the Central and South American markets. While this surge in popularity may exhibit some level of retraction post-hype in regions like North America and Pan-Europe, the overall trend suggests a solidifying presence of South Korean media on the world stage, challenging established cultural dominions and reflecting the dynamism of global VOD content consumption.
@@
It is noteworthy to mention Turkiye’s unique position in global media consumption. Turkiye occupies the most ambiguous position among the three similar groups. Data suggests that Turkiye consumes American content from the perspective of the Middle East and Africa group, whereas it consumes South Korean content from the perspective of the North America and Pan-Europe group. Such results suggests that Turkiye’s content preference align variably with different cultural groups depending on the content’s origin.
@@
The consumption behaviors of users on unprecedented global VOD content are a direct reflection of the anthropological trajectory of the world. Raymond Williams explored the gradual progression of cultural revolution in his pivotal work “Television: Technology and Cultural Form” published during the height of the mass media era in 1974. Now, precisely fifty years later in 2024, the predominant consumption of VOD content within the extensive historical and cultural trajectory sends a thrill through researchers. The consumption of global VOD content seems to manifest what Williams might call a ‘long future.’ Yet, it remains to be observed whether the recent success of unique cases such as South Korea, with hits like Squid Game and Pachinko, is an accidental ‘outcome’ or will prove to be a sustained ‘factor’ in the future.
4.2 Discussion
The current findings lead us to realize that VOD provides two significant implications that set it apart from traditional television popular culture. First is regarding the geographic landscape of VOD popular culture. Unlike the 20th-century television that defined a common culture within national borders, VOD culture unfolds over geo-linguistic and regional blocks. Unlike traditional broadcasting, which was bounded by national borders, Netflix represents a transnational, transcultural service that transcends multiple countries and unites them on a platform[21]. That is, VOD platforms have created a new type of common culture that goes beyond traditional geographic and linguistic boundaries. Thus, strictly speaking, VOD services could contribute the popular culture of global regional audiences, reflecting their tastes and, more broadly, the universal values shared among them[5, 14]. It appears that the boundaries of VOD content’s regional blocs are established at the intersection of commonality and universality. VOD services pioneer global landscapes of human interaction, media, and technology, extending beyond the realms of finance and commodities.
The global patterns we see in how people use VOD services point to a shift in the way we understand capitalism. It’s no longer just about a competitive market with many channels, as we saw with postmodern capitalism. Now, we’re entering an era of post-postmodern capitalism, where even though people across different regions and languages consume content differently, there’s a trend towards creating a shared global culture[24]. Post-postmodern culture is all about the richness and diversity that come from our connected world. It’s a culture shaped by everyone getting involved, through the interconnected network of our digital age. The way we watch shows and movies on VOD platforms shows that our personal choices are part of a bigger picture, influenced by our shared history and culture. Furthermore, the way VOD services are tailored to fit what each person wants to watch reflects this new cultural landscape. Nowadays, consuming VOD series from other countries has become a familiar landscape. Just as television consistently transcended the boundaries of the household[30], VOD services, now crossing even national borders, have evolved into a universal culture. This culture reflects the aesthetic preferences of individual viewers and fosters cultural connections with anyone worldwide who shares such cultural ties.
The second point is about the new culture shaped by VOD users worldwide. This universal culture is a complex, self-organized force created by the global participating public, reflecting the political economy unique to VOD platforms that capitalizes on viewers as a resource. Instead of being seen merely as targets for advertising, as was common in past media models, VOD users are now valued for their active engagement and the data they generate, such as their viewing habits and subscription payments[19]. This shift has led to the formation of distinct viewer groups or clusters, each with its own patterns of influence and preferences. This idea moves away from the outdated view of audiences simply as commodities[28] and towards recognizing them as resources within the VOD ecosystem. We should now talk about ‘audience resources.’ The global sourcing of finance, along with the diverse and self-directed behaviors of VOD users, is creating new dynamics in media consumption. The intricate processes of VOD consumption and the activity of viewers as resources are two sides of the same coin within this framework.
Central concepts to complexity theory, such as patterns and self-organization, offer significant conceptual and theoretical resources to platform media, where users act as nodes in a network. As networks perform self-organization, the current research findings can offer insights not only into VOD content consumption leadership within clusters but also into exploring genre similarity within them. The leadership exhibited by certain producing or consuming countries indicates a high probability that genres favored by these countries will be universally consumed within the cluster.
From this perspective, a complex systems approach is both distinctive feature and a limitation of this study. The three similar consumption groups identified here are not static; they can change their patterns at any time depending on the activities of global VOD content producers and the consumption behaviors of global viewers. The consumption of specific countries or genres, and the emotions they invoke, emerges through the interactions of individuals participating as nodes in the vast platform network. Therefore, a media complexity approach to global VOD service research is essential. It is also necessary to consider that the tendency for regional bloc formation could be influenced by local opt-out policies, where VOD services locally differentiate content based on region and country. The cultural power of VOD programming as an independent variable should not be underestimated. Nevertheless, while bloc formation can vary, the dynamism of countries with specific historical and cultural backgrounds appears to transcend such policies. Further research in this area is anticipated.
5 Conclusion
In conclusion, our study delves into the transformative realm of global VOD media consumption, revealing patterns that reflect the complex interplay of cultural, geographical, and linguistic factors. We have observed that VOD platforms, transcending traditional media boundaries, are fostering a novel, self-organized global culture, unconfined by national borders. This new era of post-postmodern capitalism heralds the rise of a so called ‘universal culture’ shaped by a networked audience whose tastes and preferences are as varied as they are interconnected[29]. When we refer to cultural universals—patterns, traits, or elements found commonly across different cultures[10]—universal culture signifies those aspects of life that are present across all human cultures, regardless of race, nationality, or region.
The leadership in content production and consumption that countries like the United States and South Korea have demonstrated suggests the emergence of new forms of cultural influence within the global landscape. Notably, the impact of significant cultural phenomena, such as South Korea’s Squid Game, challenges and redefines traditional media consumption patterns.
This study highlights the importance of considering VOD users as dynamic nodes within a vast network, where each node contributes to a collective culture that is continuously evolving. The implications of this evolution are profound, suggesting a shift away from the audience as a mere commodity towards recognizing viewers as active contributors to the media ecosystem.
The findings underscore the need for an ongoing evaluation of how global VOD content production and consumption may continue to reshape the cultural landscape of societies worldwide. As the global media landscape evolves, so too will the cultural narratives that emerge from this interconnected space. The notion of the ‘long future’ reminds us that the shifts we are witnessing are part of a larger historical continuum, carrying the potential to influence the trajectory of global culture for years to come.
6 List of Abbreviations and Symbols
Symbol/Abbreviation Meaning
NS Netflix Show
NSC Netflix Show Consumption
$N_{day}$ Number of data collection days (822 days)
$N_{coun}$ Number of countries in data collection (71 countries)
$N_{cot}$ Number of contents in data collection (1769 contents)
$g_{A}^{d,c}$ The ranking of content $c$ on day $d$ in country $A$
$G_{A}$ The set of $g_{A}^{d,c}$
$s_{A}^{d,c}$ The score of content $c$ on day $d$ in country $A$
$dd_{AB,d}$ Daily distance between country $A$ and $B$ on day $d$
$D_{A,B}$ The NSC distance between country $A$ and $B$
$W_{A,B}$ The weight between country $A$ and $B$
Table 1: List of abbreviations and symbols used in the manuscript.
7 Declarations
  • Ethics approval and consent to participate
    Not applicable
  • Consent for publication
    Not applicable
  • Availability of data and materials
    The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
  • Competing interests
    The authors declare no competing interests.
  • Funding
    Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2022S1A5A2A03051182).
  • Authors’ contributions
    N.L., J.L., and H.C.J. conceived and planned the project. N.L. conducted the computations, simulations, and generated the figures, with input from all authors. All authors contributed to writing the manuscript.
  • Acknowledgements

### v3

@@ 3.2 The role of key content-producing countries in similar consumption grouping
Figure 3: Similar consumption groups by the leading content-producing countries. (a)Similar consumption groups by US-produced content, (b)Similar consumption groups by South Korea-produced content, (c)Similar consumption groups by Colombia-produced content
Then, to what extent do the key content-producing countries influence the grouping into similar consumption clusters? To examine the extent to which major content-producing countries are associated with the formation of consumption clusters, we conducted the Louvain algorithm process again, but this time we only included the consumption data from the leading content-producing country within each of the three identified consumption groups. Networks were constructed from 528 pieces of United States content, 203 pieces of South Korean content, and 27 pieces of Colombian content.
When forming similar consumption clusters based on the leading countries’ content, most countries consistently belonged to the same group across 1,000 runs of the Louvain algorithm, similar to the previous results. A few countries, however, were classified into different groups in some runs. Even so, every country was assigned to one group at least 700 times out of 1,000, and Fig. 3 shows the groups to which countries most frequently belonged.
@@
Finally, when the content consumption patterns are analyzed with Colombian content as the focus, we observed two main cluster groups, a Hispanic Latin America and Spain group and other remaining groups. That is, when looking from the perspective of Colombia, the division of groups appears to be heavily influenced by language, specifically whether the countries are Spanish-speaking or not. This reflects the cultural and linguistic influence on media consumption within the VOD market.
@@
In sum, US content is similarly consumed in Asia, Central and South America, as well as the Middle East and Africa. Specifically, Spain is associated with the North America and Pan-Europe consumption group when it comes to US content, while Turkiye is associated with the Middle East and Africa. This indicates that audiences in Spain and Turkiye share consumption patterns of US content with North America and Pan-Europe, and with the Middle East and Africa, respectively. South Korean content consumption shows patterns that closely resemble the three consumption clusters previously identified. However, countries such as Argentina, Spain, and Turkiye display viewing patterns of South Korean content that are more similar to those of North America and Pan-Europe. Colombian content is mostly consumed within the Spanish-speaking cultural sphere, including Latin America’s central western regions and Spain in Europe, while other countries are grouped into a different cultural sphere. Overall, this analysis suggests that content from major producing countries is associated with distinct global consumption patterns, which may converge or diverge depending on cultural and linguistic ties. The way content from these countries is consumed may therefore contribute to shaping or reflecting cultural spheres and clusters within the VOD space.
3.3 Trends in content-producing countries
@@
Figure 4: Trends in the rankings of US, South Korean, and Colombian content by similar consumption groups. The red triangle represents the scores for American content, the green circle represents South Korean content, and the blue star represents Colombian content. (a)The Asia and Middle East Group, (b)The North America and Pan-Europe Group, (c)The Central and South America Group.
@@
Based on the original analysis, which defined three similar consumption groups by considering all content, we further examined how the content rankings of leading countries within the VOD similar consumption groups, such as the US, South Korea, and Colombia, have changed over the data period. This aims to assess whether the previously observed content performance reflects random variation or consistent patterns. Figure 4 represents graphs that illustrate the quarterly ranking scores of content from the United States, South Korea, and Colombia within the three established similar consumption groups.
A notable peak in the preference for South Korean content appears in the fourth quarter of 2021 and has continued to grow steadily thereafter. This pattern may be partially associated with the global popularity of Korean series released during that period, including Squid Game, which became a worldwide phenomenon. In the North America and Pan-Europe group, while American content remains strongly preferred, there is a noticeable spike in interest for South Korean content in the fourth quarter of 2021, which coincides with the release period of Squid Game, followed by a return to prior levels. In the Central and South America group, South Korean content consumption, which was previously minimal, shows a noticeable increase in the fourth quarter of 2021—a period that, once again, overlaps with the global release of Squid Game—and remains at a higher level thereafter.
The global response to South Korean content shows a noticeable change after 2021, during the same period as the release of Squid Game, which may have contributed to the shift.
4 Summary and Concluding Remarks
@@
Our findings show that global countries are categorized into three groups with geographic and linguistic connections based on global VOD consumption distance, the ‘North America and Pan-Europe group’, the ‘Asia and Middle East group’, and the ‘Central and South America group’. In these three groups, the ‘North America and Pan-Europe group’ is most influenced by the United States, the ‘Asia and Middle East group’ by South Korea, and the ‘Central and South America group’ by the United States and Colombia.
Looking specifically at the top five producing countries, the United States shows the highest share of production scores in all three groups, followed by Spain and South Korea, which contribute at intermediate levels. Other countries also stand out depending on their economic and cultural presence within the respective groups. This pattern partially aligns with the fact that Netflix’s original investment in Europe is made selectively according to existing cultural power[24].
It highlights that the United States accounts for the largest share of content consumption, the distinctive geographical and linguistic reach of Spain, and the role of countries with notable production shares within each group. South Korea is an exceptional case that cannot be explained solely by geography or language within this framework, warranting further examination.
The rise in South Korean content popularity observed in the fourth quarter of 2021 coincides with the release of the highly popular series Squid Game. After this period, the preference for South Korean content consistently increased in the ‘Asia and Middle East group’ and the ‘Central and South America group’. In contrast, in the ‘North America and Pan-Europe group’, the preference for South Korean content subsided relatively quickly, likely reflecting the strong presence of US content in that region. These observations suggest that the release of widely recognized content such as Squid Game may be associated with shifts in global consumption patterns, even in regions with established viewing tendencies. The increased attention to South Korean content following Squid Game indicates a notable change in its global visibility and consumer appeal, with especially strong gains in Asia and meaningful presence in the Central and South American markets. While this surge in popularity appears to diminish more rapidly in regions such as North America and Pan-Europe, the broader trend points to an expanding role of South Korean media within global VOD consumption, highlighting the fluid and dynamic nature of content preferences across cultural spheres.
@@
It is noteworthy to mention Turkiye’s unique position in global media consumption. Turkiye occupies the most ambiguous position among the three similar groups. Data suggests that Turkiye consumes American content from the perspective of the Middle East and Africa group, whereas it consumes South Korean content from the perspective of the North America and Pan-Europe group. Such results suggest that Turkiye’s content preferences align variably with different cultural groups depending on the content’s origin.
Viewing users’ engagement with global VOD platforms within a broader anthropological and cultural trajectory, Raymond Williams explored the gradual progression of cultural revolution in his pivotal work “Television: Technology and Cultural Form” published during the height of the mass media era in 1974. Today, about fifty years later, the widespread consumption of VOD content within this long historical and cultural trajectory offers new opportunities for researchers to explore.
@@
The consumption of global VOD content seems to manifest what Williams might call a ‘long future’, a notion that emphasizes the gradual, cumulative processes of cultural transformation rather than a fixed or distant endpoint. Yet, it remains to be observed whether the recent success of unique cases such as South Korea, with hits like Squid Game and Pachinko, is an accidental ‘outcome’ or will prove to be a sustained ‘factor’ in the future.
In conclusion, our study examines global VOD media consumption and identifies clear patterns shaped by cultural, geographical, and linguistic associations[25]. Using similarity networks, we find that countries consistently cluster into three major groups, reflecting both regional proximity and shared cultural characteristics. The analysis also shows that content from major producing countries, such as the United States, South Korea, and Colombia, is associated with distinctive global consumption patterns, which may converge or diverge depending on historical and linguistic ties[26]. These results highlight how VOD platforms, while transcending national boundaries, still reflect underlying cultural structures and offer a framework for understanding the ways in which global audiences organize around shared viewing preferences.
Future research should examine whether current shifts, such as the growing popularity of non-Western content, will become lasting features of the global media landscape or prove to be short-lived. In addition, it would be an intriguing topic to explore not only the contemporaneous relationships among countries but also the temporal dynamics of content spread across nations[27]. VOD media, positioned at the intersection of technology, culture, and audience choice, will remain a key area for studying how global and local cultural forces interact over time.
Appendix Appendix A Validation of the Method
In this appendix, we examine the validity of our methodology and test whether our results are robust under different assumptions. As explained in the main text, our analysis is based on Netflix Top-10 show rankings for each country from July 2020 to September 2022. The dataset contains $N_{\text{cont}}=1769$ unique contents recorded across $N_{\text{day}}=822$ days. Rankings were converted into scores, with 10 points assigned to 1st place, 9 points to 2nd place, and so on down to 1 point for 10th place, while contents outside the top-10 list received 0 points. Each country was then represented as a point in a high-dimensional space of $822\times 1769=1,454,118$ dimensions, with coordinates determined by these scores. Distances between countries were calculated using the Euclidean distance, and then converted into weights $W_{\!AB}$ using Eq. (7) in the main text.
This method provides a systematic and interpretable way to measure similarities in consumption patterns. However, two potential limitations require further validation. First, the availability of specific content can differ across countries. Second, our analysis uses only the top-10 rankings, which means we cannot distinguish between contents that were unranked and those that narrowly missed the top-10.
To address these issues, we test whether restricting the analysis to the top-10 is a valid approach and then explore alternative definitions of distance between countries. By doing so, we assess whether the similarity patterns observed in the main text remain robust under these variations.
A.1 Distribution of Content Scores
Figure A1: (a) Total scores $s_{{}_{\!T}}^{c}$ of $N_{\text{cont}}=1769$ Netflix titles arranged by rank, plotted on a log–log scale. (b) Distribution of total scores $s_{{}_{\!T}}^{c}$ on a log–log scale using bins of width 10. The vertical axis indicates the number of contents $n(s_{{}_{\!T}}^{c})$ falling within each bin.
In this section, we examine the distribution of content scores to understand the variation in popularity among Netflix shows. Since direct viewership data are not available, we use the total score $s_{{}_{\!T}}^{c}$ defined in Eq. (6) as a proxy for global popularity. This quantity sums the ranking scores of each content across all countries and days in the dataset.
Figure A1 presents two complementary views of this distribution. Panel (a) shows the total scores $s_{{}_{\!T}}^{c}$ of $N_{\text{cont}}=1769$ contents arranged by rank on a log–log scale. The most popular contents reached extremely high scores, with the top five being 1. Yo soy Betty, la fea with $62,671$ points, 2. Bridgerton with $56,828$ points, 3. Stranger Things with $50,364$ points, 4. Pablo Escobar, The Drug Lord with $49,617$ points, and 5. Squid Game with $49,527$ points. By contrast, a large portion of the catalog attracted very little attention, with 671 out of 1,769 titles scoring fewer than 100 points. This highlights the highly uneven nature of global consumption, where only a small number of titles dominate.
Panel (b) shows the distribution of $s_{{}_{\!T}}^{c}$ on a log–log scale using fixed-width binning. Each bin has a width of 10, so the first bin includes scores from 1 to 10, the second from 11 to 20, the third from 21 to 30, and so on. The vertical axis indicates the number of contents $n(s_{{}_{\!T}}^{c})$ in each bin. The figure clearly demonstrates a heavy-tailed pattern, where a few contents accumulate very high scores while most remain at extremely low levels.
Together, these results highlight the highly uneven distribution of global content popularity, where a small number of blockbuster shows attract overwhelming attention compared to the majority of other titles. Hence, this analysis indicates that the top-10 lists in each country may capture enough information for the study.
A.2 Robustness Analysis on Similarity Metrics
Figure A2: Comparison of the weights $W_{\!AB}$ between all $N_{\text{coun}}(N_{\text{coun}}-1)/2=2485$ pairs of countries calculated via different methods. Across all subfigures, the x-axis represents weights computed by the original method using Eq. (7) in the main text. The y-axis in each subfigure corresponds to weights recalculated using alternative methods, (a) considering only top 5 content instead of top 10 content (alt-method 1), (b) assigning a distance of 5.5 when a title was outside the top-10 in both countries (alt-method 2), (c) including only content that appeared in the top 10 list of both countries (alt-method 3), and (d) computing cosine similarity between country vectors instead of Euclidean distance (alt-method 4). All panels show strong correlations, supporting the robustness of the original method.
Here, we consider different ways of defining similarity weights between two countries and test whether the observed patterns remain consistent under these alternatives. For the four alternative methods, we present scatter plots in which the $x$-axis represents the weights computed by the original method using Eq. (7) in the main text, and the $y$-axis corresponds to the weights recalculated using the alternative methods. To quantify the consistency, we calculate the Pearson correlation coefficient $R$, which measures the strength of a linear relationship between two variables and is defined as
$R$ $=$ $\frac{\sum_{i=1}^{N}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}\,\sqrt{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}},$ (10)
where $x_{i}$ and $y_{i}$ are paired data points, and $\bar{x}$ and $\bar{y}$ are their respective means. We also report the coefficient of determination $R^{2}$, which measures how much of the variation in one variable can be explained by the other. It is defined as
$R^{2}$ $=$ $1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}},$ (11)
where $\text{SS}_{\text{res}}=\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$ is the residual sum of squares, and $\text{SS}_{\text{tot}}=\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}$ is the total sum of squares. Here, $\hat{y}_{i}$ denotes the fitted value from the regression and $\bar{y}$ is the mean of the observed data.
We begin by testing whether using only the top 10 rankings is a valid approach for calculating distances between countries. Since data beyond the top 10 is not available, all lower-ranked content is treated the same, which may overlook meaningful differences in viewer preferences. To assess the impact of this limitation, we calculate alternative weights $W_{\!AB}^{(5)}$ using only the top 5 rankings (referred to as alternative method 1) and examine whether the resulting weight ordering depends on the ranking depth. In this version, the number of content items is reduced from 1,769 to 1,314. Scores are assigned from 5 to 1 for ranks 1st through 5th, and 0 for all other ranks. Figure A2(a) compares these similarity weights with those from the original top-10-based method. The high correlation between the two sets of weights, with Pearson correlation coefficient $R=0.992$ and the coefficient of determination $R^{2}=0.985$, suggests that our main results are not strongly affected by the ranking depth, supporting the robustness of our approach across different cutoff choices.
Next, we examine the implication of assigning a uniform score of 0 to content not in the top 10. Such an assignment nullifies its contribution to distance even if the unranked content differs substantially across countries. To approximate possible differences among unranked items, we instead assign a notional distance value of 5.5 when a title is ranked 11th or lower in both countries. This reflects the average value of scores within the top 10, introducing a uniform middle-ground contribution to the distance measure. Yet, the selection of 5.5 as the constant value is somewhat arbitrary but is used to illustrate that incorporating a specific, finite value for ranks beyond the top 10 does not markedly influence our findings. Figure A2(b) shows the comparison of weights under this definition (referred to as alternative method 2) and the original definition. A strong correlation is again observed, with $R=0.983$ and $R^{2}=0.967$ suggesting that the particular treatment of unranked content does not substantially alter the global similarity patterns.
The timing of content releases and the availability of specific shows vary across countries, which could influence our weight calculations. To assess the robustness of our method to such differences, we conducted an alternative analysis using only content that simultaneously appeared in the top 10 lists of both countries on the same day. By focusing exclusively on shared content, we eliminate potential biases caused by differences in availability. Figure A2(c) compares the resulting weights from this method (referred to as alternative method 3) with those obtained using the original definition. Once again, we observe a quite strong positive correlation between the two sets of weights, with Pearson correlation coefficient $R=0.926$ and $R^{2}=0.858$ indicating that our similarity measure remains consistent even when restricted to commonly available content.
Finally, we examine whether our results depend on the choice of distance metric itself. All previous calculations used Euclidean distance between country vectors. As an alternative, we compute cosine similarity between country vectors, defined as the cosine of the angle between the two high-dimensional vectors. This similarity is often used in information retrieval and recommendation systems because it focuses on directional similarity, regardless of magnitude. We then transform the cosine similarity values into a corresponding distance measure and calculate the weights $W_{\!AB}^{(5)}$ accordingly (alternative method 4). Figure A2(d) compares these weights with those from the original Euclidean-based method. We again observe a high correlation ($R=0.991$ and $R^{2}=0.982$), indicating that the overall similarity patterns among countries are not sensitive to the choice between Euclidean and cosine distance. Together, these findings confirm the robustness of our approach under a range of plausible variations.
Acknowledgments